OpenStudy (anonymous):

find the slope (-2,6)(4,3)

OpenStudy (anonymous):

Can you plug the numbers into this formula: y=mx+b?

OpenStudy (radar):

When two points are given you can find slope by: m=(6-3)/(-2-4) or m=(y2=y1)/(x2-x1) in this case m=3/-6=-(1/2)

OpenStudy (radar):

m=(y2-y1)/x2-x1) \[m =\Delta y/\Delta x\]

OpenStudy (radar):

Was that helpful?